Abigail is putting juice into cone - shaped containers to make popsicle's. She has approximately one quarter of a gallon, or 58 cubic inches, of juice to make the popsicle's with. Each container has a height of 3 inches.


Use the information to complete the table indicating the maximum number of popsicle's that Abigail could make with 58 cubic inches of juice.

Radius of the
Container
0.5 inch
1 inch
1.5 inch
What is the maximum number of popsicle's per container?

Answer by mananth(16946)


Abigail is putting juice into cone - shaped containers to make popsicle's. She has approximately one quarter of a gallon, or 58 cubic inches, of juice to make the popsicle's with. Each container has a height of 3 inches.

Use the information to complete the table indicating the maximum number of popsicle's that Abigail could make with 58 cubic inches of juice.
Radius of the
Container
0.5 inch
1 inch
1.5 inch
What is the maximum number of popsicle's of each size
Volume of container = 1/3 *pi*r^2*h
radius =0.5 in
height = 3 inches
Quantity she has is 58 in^3
V=(1/3)*pi*(0.5)^2*3=0.785 in^3 per popsicle
Number of popsicles = 58/0.785 =73.88
So 73 popsicles from the quantity she has with her
repeat for other sizes